A total of $10,000 is invested in two funds paying 5% and 7% annual interest. The combined annual interest is $630. How much of the $10,000 is invested in each fund?
Answer by stanbon(75887)

A total of $10,000 is invested in two funds paying 5% and 7% annual interest. The combined annual interest is $630. How much of the $10,000 is invested in each fund?
----
Equation:
interest + interest = 630
0.05x + 0.07(10,000-x) = 630
----
Multiply thru by 100 and solve for "x":
5x + 7*10,000 - 7x = 63000
-2x = -7000
x = $3500 (amt. invested at 5%)
10,000-x = $6500 (amt. invested at 7%)
==========================================
Cheers,
Stan H.
===============